# Harnessing Vision-Language Models for Time Series Anomaly Detection

### v1

@@ 4.1 Experimental Setup
Benchmark Datasets and Baseline Methods. For unsupervised anomaly detection, we follow the standard evaluation protocol to test ViT4TS and VLM4TS on 11 widely used benchmark datasets in time series anomaly detection research [1, 2], spanning various domains from sensor data like astronomy sensory, web monitoring data like production traffic, to web metric data like volume of Twitter mentions, to evaluate the models’ generalizability and adaptability. For baseline models, we compare our method against several anomaly detection approaches, from statistical baselines like ARIMA [11], to deep learning models currently considered state-of-the-art, including a forecasting- based LSTM model (LSTM-DT) [2], (variational) reconstruction-based models like LSTM-AE [12], VAE [41], and TadGAN [14], hybrid models like AER [13], transformer-based models like Anomaly Transformer (ATrans) [18], and pre-trained time series foundation models like UniTS [21] and TimesFM [25], as well as LLM-based approaches such as prompt-based detectors (SigLLM-PG on GPT, SigLLM-PM on Mistral) and LLM prediction-based models SIGLLM-D [5]. We also report TAMA [7], a naive rolling-window VLM prompting baseline.
@@
Evaluation. We follow the standard unsupervised TSAD protocol [42, 43]. For any method that produces continuous anomaly scores, we first smooth the raw outputs with an exponentially weighted moving average (EWMA). We then apply a Gaussian-based threshold of the form $\mu+k\sigma$, where $\mu$ and $\sigma$ are the mean and standard deviation of the smoothed scores, sweeping $k$ to compute the unweighted contextual F1 score [2, 13, 14]. Max F1 score (F1-max) results appear in the main text; full F1 results are provided in the Appendix B.6. Methods yielding only binary labels (SigLLM-PM, SigLLM-PG) are evaluated at their default settings and reported by raw F1. All experiments run on an NVIDIA V100 GPU if not otherwise specified. Baselines are implemented via the Orion framework [43] when available; otherwise, we follow each original implementation’s setup and prompt.
Setup. For the primary experiment, ViT4TS uses the OpenCLIP [44] implementation of CLIP with the ViT-B/16 backbone, pretrained on the LAION-400M dataset [36]. For multi-scale embedding extraction, we perform average pooling over the patch grid using kernel sizes of $2\times 2$ and $3\times 3$. For the verification stage, VLM4TS employs GPT-4o (API version 2024-08-06) via OpenAI’s API.
@@
Table 1: Detection performance (F1-max) of ViT4TS and VLM4TS versus trained-from-scratch and time-series-pretrained baselines on benchmark datasets. Each entry reports the maximum F1 score across all evaluated thresholds; the best score is shown in bold, and the second-best is underlined. Full F1 results and elapsed (wall-clock) time comparison are provided in Appendix B.3.
@@
Trained From Scratch ARIMA 0.387 0.263 0.500 0.344 0.179 0.585 0.750 0.650 0.771 0.502 0.336 0.479$\pm$0.187
AER 0.338 0.244 0.518 0.404 0.178 0.553 0.753 0.618 0.866 0.711 0.614 0.527$\pm$0.207
TadGAN 0.338 0.196 0.385 0.421 0.205 0.612 0.593 0.492 0.667 0.135 0.109 0.378$\pm$0.189
LSTM-DT 0.368 0.273 0.444 0.451 0.190 0.615 0.724 0.639 0.877 0.704 0.534 0.529$\pm$0.200
LSTM-AE 0.231 0.244 0.400 0.416 0.232 0.487 0.673 0.583 0.853 0.584 0.201 0.446$\pm$0.203
ATrans 0.262 0.168 0.200 0.365 0.147 0.454 0.567 0.263 0.554 0.437 0.394 0.346$\pm$0.142
VAE 0.000 0.248 0.345 0.323 0.237 0.515 0.695 0.557 0.845 0.524 0.189 0.407$\pm$0.234
Time-series Pretrained TimesFM 0.234 0.243 0.400 0.467 0.198 0.564 0.686 0.554 0.694 0.120 0.107 0.388$\pm$0.209
UniTS 0.182 0.246 0.326 0.479 0.167 0.561 0.723 0.605 0.760 0.126 0.110 0.390$\pm$0.233
Ours ViT4TS 0.545 0.400 0.615 0.615 0.597 0.543 0.726 0.614 0.892 0.614 0.565 0.612$\pm$0.116
VLM4TS 0.714 0.488 0.727 0.632 0.686 0.619 0.773 0.733 0.901 0.497 0.474 0.659$\pm$0.127
@@
F1-max Tokens Time F1-max Tokens Time F1-max Tokens Time F1-max Tokens Time
LLM-based SIGLLM-D 0.353 11153 83.77 0.232 13157 89.50 0.393 21508 62.20 0.326 15273 78.49
SigLLM-PM 0.206 14191 613.28 0.157 45050 2107.96 0.276 45082 984.45 0.213 34774 1235.23
SigLLM-PG 0.162 25006 2258.78 0.080 78187 2614.87 0.143 83207 2852.52 0.128 62133 2575.39
VLM-based TAMA 0.513 40009 110.47 0.631 31563 83.49 0.616 27324 68.99 0.587 32965 87.65
@@ 4.2 Main Results
Overall Detection Performance. Table 1 reports the F1-max scores versus trained-from-scratch and time-series-pretrained baselines on all 11 benchmark datasets. Our two-stage VLM4TS framework achieves the highest average F1-max, outperforming competing methods on 9 out of 11 datasets, achieving a 24.6 % improvement in average F1-max score over the second-best baseline LSTM-DT. These results underscore that purely vision-driven screening, when coupled with powerful VLM reasoning, can exceed state-of-the-art time series-based TSAD models in most cases. Remarkably, even the first-stage visual screening module ViT4TS—without any time series–specific training—ranks second overall, securing the top-two position on 6 datasets, showing the effectiveness of detecting time series anomalies from a visual perspective. In Table 2, we further compare against existing language model-based baseline methods, where VLM4TS also consistently outperform baselines by a large margin, achieving 13.3% performance gain over the VLM prompting-based method TAMA and ×$\times 2$ improvement over the strongest LLM-based method SigLLM-D. Moreover, we observe that both VLM-based methods (ours and TAMA) substantially surpass LLM prompting frameworks such as SigLLM-PM and SigLLM-PG, confirming that casting time series as images unlocks more effective anomaly detection than text-only representations (Figure 1-(a)).
@@
Task-Specific Detection Performance. Compared with time-series-based methods, VLM4TS delivers its strongest gains on real-world datasets dominated by contextual anomalies (e.g., the NAB dataset group), where it outperforms all baselines by a large margin. Its advantage stems from the ability to localize anomalies in a 2-D representation and then apply global-context verification to boost precision without sacrificing recall. See Appendix B.4 for detailed results and related discussion on precision and recall. In contrast, on a few synthetic benchmarks A3 and A4—where anomalies are densely and synthetically injected (average anomalies per time series: A3 = 9.39, A4 = 8.37 versus A1 = 2.66, A2 = 2.00)—forecasting-based methods (e.g., AER and LSTM-DT) excel by design. Under our rarity assumption, VLM4TS adopts a conservative thresholding strategy that overly reduces recall in these high-density settings (Appendix B.4). However, we claim that this conservatism aligns with practical industrial scenarios, where minimizing false positives is critical to operational efficiency. Compared with language-based methods like rolling-window VLM prompting method TAMA, VLM4TS delivers the strongest gains in datasets that demand long-range context, most notably the NAB dataset group (26.5%). This advantage arises from our two-stage design: ViT4TS supplies high-resolution, localized proposals, allowing VLM4TS to verify and refine anomalies over a much larger temporal horizon without losing localization accuracy.
@@
Computational Time and Token Cost. Table 2 compares F1-max score, per-series elapsed time, and token consumption for VLM4TS versus language model-based baseline methods. We also report per-series elapsed time comparison with time-series-based method in Appendix B.3. Due to our two-stage design, VLM4TS requires much less token cost, reducing token usage by an average of 30× compared to pure LLM- and VLM-based detectors that encode numerical data into text or image for every rolling window. When compared against time series–only approaches—both pretrained foundation models and models trained from scratch, VLM4TS achieves comparable end-to-end runtime while achieving better overall detection accuracy. We also evaluate CPU-only operation of VLM4TS (with VLM running on the API) with the default ViT-B/16 backbone. On moderate-length series spanning thousands of time points, ViT4TS screening completes in seconds, demonstrating feasibility for large-scale deployment without GPU acceleration and with low token cost.
@@
4.3 Additional Analysis
$\bullet$ Multi-scale embedding: ablating multi-scale embedding extraction reduces performance, especially on NASA dataset group, which can lead to a 8.35% drop in F1-max, as multi-scale embeddings enhance detection to extended contextual anomalies that cannot be captured by patch-level comparison, particularly in domains like spacecraft telemetry.
@@ 4.3 Additional Analysis
VLM Ablation. To assess the effect of the VLM model for the verification stage, we replace GPT-4o with alternatives such as Gemini and Claude and present the full results in Appendix B.2. We observe that some models take a conservative verification, confirming only the strongest candidates and thus yielding higher precision but lower recall, while others adopt a more aggressive approach, boosting recall while incurring lower precision.
Figure 5: Qualitative comparison of anomaly detection on MSL C-1 (left) and SMAP A-4 (right). "VLM-Long” refers to VLM prompting on the full-series image (ablation without ViT4TS screening); “VLM-Short” (TAMA) refers to VLM prompting on the rolling-windows image. Only representative segments are shown for clarity.
Embedding and Anomaly Map Analysis To demonstrate ViT4TS’s ability to separate anomalous regions from normal patterns, we visualize patch-level embeddings and compute a corresponding anomaly score heatmap for a representative time series anomaly. See Appendix B.5 for details. After projecting the patch embeddings onto two dimensions using t-SNE, we observe anomaly-covering patches form a clearly distinct cluster from normal patches, showing that the pretrained Vision Transformer captures local shape deviations. In addition, the cross-patch dissimilarity heatmap highlights the same anomaly as a contiguous bright band, confirming precise localization on the 2-D anomaly map before mapping back to 1-D anomaly score.
@@
Qualitative Results. Figure 5 illustrates anomaly localization on two NASA telemetry signals. In both cases, VLM4TS delivers the most precise detections: it refines ViT4TS’s candidate proposals by comparing each candidate against the full-series context, correctly isolating the true anomalous intervals. By contrast, the “VLM-Long” ablation—prompting the VLM on the entire series without prior screening—either misaligns its anomaly boundaries (MSL C-1) or fails to flag the event entirely (SMAP A-4). Meanwhile, “VLM-Short” (TAMA), which prompts on every rolling window, generates numerous false positives due to its narrow context and incurs prohibitively high token usage.
5 Conclusion
@@
We introduce a novel two-stage framework for unsupervised time-series anomaly detection that reframes 1-D time series as 2-D visual inputs and leverages pretrained vision and vision–language models without any in-domain training. In the first stage, ViT4TS applies multi-scale cross-patch comparisons to generate anomaly candidates in high resolution, and in the second stage, VLM4TS verifies these candidate proposals via temporal reasoning with long context. Our method achieves state-of-the-art performance on diverse benchmarks while using much fewer tokens.
Limitation and Future Work. While our results demonstrate VLMs’ promise for unsupervised anomaly detection, our current prompting strategy remains simple. Future work could explore advanced in-context reasoning methods, such as chain-of-thought or retrieval-augmented prompts, to further enhance temporal inference. Additionally, we have focused on univariate series; extending to multivariate data will require new visualization and encoding schemes (for example, overlaying multiple signals) to capture inter-series relationships. We leave these explorations for future work.
Appendix A Experimental Details
A.1 Benchmark Datasets
@@
We use 11 datasets spanning various domains to evaluate the models’ generalizability and adaptability. The National Aeronautics and Space Administration (NASA) provided two spacecraft telemetry datasets  [^1] NASA data: https://github.com/khundman/telemanom/: Soil Moisture Active Passage (SMAP) and Mars Science Laboratory (MSL) acquired from a satellite and a rover, respectively [2]. Each numeric measurement in the target channel is accompanied by one-hot encoded information about commands sent or received by specific spacecraft modules in a given time window. The Yahoo Webscope Program provided the S5 datasets  [^2] YAHOO data: https://webscope.sandbox.yahoo.com/catalog.php?datatype=s&did=70 consisting of one set of real production traffic to Yahoo properties (A1) and three synthetic datasets (A2, A3, A4) with varying trends, noise, and pre-specified or random seasonality. The A2 and A3 datasets only contain outliers inserted at random positions, while A4 has outliers and change points. The Numenta Anomaly Benchmark (NAB) provided several datasets  [^3] NAB data: https://github.com/numenta/NAB from various domains: artificialWithAnomaly (Art), realAdExchange (AdEx), realAWSCloudwatch (AWS), realTraffic (Traffic), realTweets (Tweets). NAB is challenging due to its variety and the requirement for detectors to work across domains. Results on NAB can assess how well our approach handles completely different kinds of time series in one unified model. Similar to [13], Table 4 summarizes the basic information of each dataset. It differentiates between real and synthetic datasets and provides the number of anomalies for each dataset. Each anomaly is classified as either point or collective, depending on the length of the anomaly. Lastly, the total number of anomalous and overall data points are provided for each dataset.
@@
A.2 Baseline Methods and Evaluation Procedure
@@
• ARIMA [11]: autoregressive integrated moving average forecasting model; learns autocorrelations to predict future values and uses point-wise forecast errors as anomaly scores.
@@
• LSTM-DT [2]: two-layer LSTM for one-step ahead prediction; uses residuals as anomaly scores.
@@
• LSTM-AE [12]: LSTM autoencoder over sliding windows; use reconstruction error as anomaly scores.
@@
• VAE [41]: variational autoencoder with LSTM encoder/decoder; use reconstruction likelihood and point-wise error to construct anomaly scores.
@@
• AER [13]: joint autoencoder and bidirectional LSTM predictor; optimizes reconstruction and forward/backward prediction, combining DTW-based and point-wise errors for anomaly scores.
@@
• TadGAN [14]: GAN with LSTM generator and critic; computes anomaly score via dynamic-time-warping, point-wise, and area-difference errors between real and generated series.
@@
• ATrans (Anomaly Transformer) [18]: transformer with anomaly-attention mechanism measuring association discrepancies among adjacent points as anomaly scores.
@@
• UniTS [21]: unified multi-task transformer pretrained on diverse series; tokenizes forecasting, imputation, and anomaly tasks for cross-domain transfer. Using forecasting error as anomaly scores.
@@
• TimesFM [25]: decoder-style attention model pretrained on patched time-series corpus; using forecasting error as anomaly scores.
@@
• SigLLM-P (SigLLM-PG on GPT, SigLLM-PM on Mistral) [5]: prompt-based detectors that ask an LLM to mark anomalous indices in textual form.
@@
• SigLLM-D [5]: detects anomalies via LLM forecasting errors—comparing next-step predictions to actual values.
@@
• TAMA [7]: rolling-window VLM prompts, rendering each window as an image and querying GPT-4o for anomaly detection, using confidence score as anomaly scores.
@@
All baselines except SigLLM and TAMA are evaluated using the Orion pipeline  [^4] Orion: https://github.com/sintel-dev/Orion. SigLLM is run via its repository  [^5] SIGLLM: https://github.com/sintel-dev/sigllm, while TAMA—which lacks a public implementation—is reproduced using the original prompts. TAMA is implemented on 20% of the series, and SigLLM-PG is on 5% of the series [5] due to high API costs.
A.3 Data Preprocessing, Postprocessing and Evaluation.
@@
Each raw series $x_{1:T}$ is first min–max normalized to $[0,1]$ and detrended by removing a least-squares linear fit; for the A4 dataset, we additionally standardize before and after the known change point. After generating anomaly scores, we apply an exponentially weighted moving average (EWMA) to smooth each score, following the unsupervised TSAD protocol of [42, 43]. We then apply Gaussian quantile thresholds $\tau_{\alpha}=\mu+k\sigma=\mu+z_{\alpha}\,\sigma$, where $\mu$ and $\sigma$ are the mean and standard deviation and $z_{\alpha}$ is the standard normal deviate at quantile $\alpha\in\{0.10,0.01,0.001\}$, and report the unweighted contextual F1 and the maximum F1 (F1-max) over these $z_{\alpha}$ values [2, 13, 14]. No interval padding is applied, ensuring that the detection performance reflects each method’s raw localization capability.
A.4 Prompt Design for VLM4TS
@@
We prompt the VLM with the full-series line plot and indicate the intervals ViT4TS marked as anomalous. Each candidate anomaly is precisely localized through the x-axis tick marks. Our goal for the VLM is to refine the output of the ViT4TS and increase the precision scores. It directs the model to discard any intervals that align with the overall trend, add missed segments exhibiting clear spikes or level shifts, and retain the exact tick-based boundaries from ViT4TS. A three-level confidence scale (1–3) supports downstream filtering. Please see the following prompting box for details.

### v2

@@ Introduction
Benchmark Datasets and Baseline Methods. For unsupervised anomaly detection, we follow the standard evaluation protocol to test ViT4TS and VLM4TS on 11 widely used benchmark datasets in time series anomaly detection research (lavin2015evaluating; hundman2018lstmdt), spanning various domains from sensor data like astronomy sensory, web monitoring data like production traffic, to web metric data like volume of Twitter mentions, to evaluate the models’ generalizability and adaptability. For baseline models, we compare our method against several anomaly detection approaches, from statistical baselines like ARIMA (pena2013arima), to deep learning models currently considered state-of-the-art, including a forecasting- based LSTM model (LSTM-DT) (hundman2018lstmdt), (variational) reconstruction-based models like LSTM-AE (malhotra2016lstmae), VAE (park2018vae), and TadGAN (geiger2020tadgan), hybrid models like AER (wong2022aer), transformer-based models like Anomaly Transformer (ATrans) (xu2021anomaly), and pre-trained time series foundation models like UniTS (gao2024units) and TimesFM (das2024decoder), as well as LLM-based approaches such as prompt-based detectors (SigLLM‐PG on GPT, SigLLM‐PM on Mistral) and LLM prediction-based models SIGLLM-D (alnegheimish2024can). We also report TAMA (zhuang2024see), a naive rolling-window VLM prompting baseline.
@@
Evaluation. For any method that produces continuous anomaly scores, we first smooth the raw outputs with an exponentially weighted moving average. We then apply a Gaussian-based threshold of the form $\mu+k\sigma$, where $\mu$ and $\sigma$ are the mean and standard deviation of the smoothed scores, sweeping $k$ to compute the unweighted contextual F1 score (geiger2020tadgan; wong2022aer). Max F1 score (F1-max) results appear in the main text; full F1 results are provided in the Appendix. Methods yielding only binary labels are evaluated at their default settings and reported by raw F1. See Appendix for the formal definition of these metrics. All experiments run on an NVIDIA V100 GPU if not otherwise specified. Baselines are implemented via the Orion framework (alnegheimish2022orion) when available; otherwise, we follow each original implementation’s setup and prompt.
vs. detection delay).
@@
Table 1: Detection performance (F1-max) of ViT4TS and VLM4TS versus trained-from-scratch and time-series-pretrained baselines on benchmark datasets. Each entry reports the maximum F1 score across all evaluated thresholds; the best score is shown in bold, and the second-best is underlined. Definition of the F1-max score, full F1 results and elapsed (wall-clock) time comparison are provided in the Appendix.
@@
Trained
From
Scratch
[…]
Time-series
Pretrained
[…]
@@
F1‐max Tokens Time F1‐max Tokens Time F1‐max Tokens Time F1‐max Tokens Time
LLM‐based SIGLLM‐D 0.353 11153 83.77 0.232 13157 89.50 0.393 21508 62.20 0.326 15273 78.49
SigLLM‐PM 0.206 14191 613.28 0.157 45050 2107.96 0.276 45082 984.45 0.213 34774 1235.23
SigLLM‐PG 0.162 25006 2258.78 0.080 78187 2614.87 0.143 83207 2852.52 0.128 62133 2575.39
VLM‐based TAMA 0.513 40009 110.47 0.631 31563 83.49 0.616 27324 68.99 0.587 32965 87.65
@@
Overall Detection Performance. Table 1 reports the F1-max scores versus trained-from-scratch and time-series-pretrained baselines on all 11 benchmark datasets. Our two-stage VLM4TS framework achieves the highest average F1-max, outperforming competing methods on 9 out of 11 datasets, achieving a 24.6 % improvement in average F1-max score over the second-best baseline LSTM-DT. These results underscore that purely vision-driven screening, when coupled with powerful VLM understanding, can exceed state-of-the-art time series-based TSAD models in most cases. Furthermore, even the first-stage visual screening module ViT4TS ranks second overall, securing the top-two position on 6 datasets, showing the effectiveness of detecting time series anomalies from a visual perspective. In Table 2, we further compare against existing language model-based baseline methods, where VLM4TS also consistently outperform baselines by a large margin, achieving 13.3% performance gain over the VLM prompting-based method TAMA and $\times 2$ improvement over the strongest LLM-based method SigLLM-D. Moreover, we observe that both VLM-based methods substantially surpass LLM prompting frameworks, confirming that casting time series as images unlocks more effective anomaly detection than text-only representations.
@@
Task-Specific Detection Performance. Compared with time-series-based methods in Table 1, VLM4TS delivers its strongest gains on real-world datasets dominated by contextual anomalies (e.g., the NAB dataset group), where it outperforms all baselines by a large margin. Its advantage stems from the ability to localize anomalies in a 2-D representation and then apply global-context verification to boost precision without sacrificing recall (see discussion on precision and recall in the Appendix). In contrast, on a few synthetic datasets, such as A3 and A4, where anomalies are densely and synthetically injected (average anomalies per time series: A3 = 9.39, A4 = 8.37 versus A1 = 2.66, A2 = 2.00), we see forecasting-based methods (e.g., AER and LSTM-DT) excelling compare to our method. Because VLM4TS assumes anomalies are rare, it adopts a more conservative behavior in high-density scenarios: visually similar, tightly packed anomaly points are interpreted as fluctuations within a broader pattern rather than as anomalies. This conservativeness results in fewer selected anomalies, leading to lower F1 scores on A3 and A4 (see Appendix for further discussion). Compared with language-based methods shown in Table 2, such as TAMA, VLM4TS delivers the strongest gains in datasets that demand long-range context, most notably the NAB dataset group (26.5%). This advantage arises from our two-stage design: ViT4TS supplies high-resolution, localized proposals, allowing VLM4TS to verify and refine anomalies over a much larger temporal horizon without losing localization accuracy.
@@
Computational Time and Token Cost. Table 2 shows the F1-max score, per-series elapsed time, and token consumption for VLM4TS and language model-based baseline methods. We also report per-series elapsed time comparison with time-series-based method in the Appendix. Due to our two-stage design, VLM4TS requires substantially less cost, reducing token usage by an average of 30× compared to pure LLM- and VLM-based detectors that encode numerical data into text or image for every rolling window. When compared against time series–only approaches—both pretrained foundation models and models trained from scratch, VLM4TS achieves comparable end-to-end runtime while achieving better overall detection accuracy. We also evaluate CPU-only operation of VLM4TS (with VLM running on the API) with the default ViT-B/16 backbone. On moderate-length series spanning thousands of time points, ViT4TS screening completes in seconds, demonstrating feasibility for large-scale deployment without GPU acceleration and with low token cost.
Additional Analysis
$\bullet$ Multi-scale embedding: ablating multi-scale embedding extraction reduces performance, especially on NASA dataset group, which can lead to a 8.35% drop in F1-max, as multi-scale embeddings enhance detection to extended contextual anomalies that cannot be captured in a patch-level, particularly in domains like spacecraft telemetry.
@@
Qualitative Results. Figure 5 illustrates anomaly localization on two NASA telemetry signals. In both cases, VLM4TS delivers the most precise detections: it refines ViT4TS’s candidate proposals by comparing each candidate against the full-series context, correctly isolating the true anomalous intervals. By contrast, the “VLM-Long” ablation (prompting the VLM on the entire series without prior screening) either misaligns its anomaly boundaries (MSL C-1) or fails to flag the event entirely (SMAP A-4). Meanwhile, “VLM-Short” (TAMA), which prompts on every rolling window, generates numerous false positives due to its narrow context and incurs prohibitively high token usage.
Figure 5: Qualitative results on MSL C-1 (top) and SMAP A-4 (bottom), illustrating how VLM4TS refines the initial local anomaly proposals from ViT4TS by selectively keeping, discarding, or adding intervals based on global temporal understanding. ”VLM-Long” refers to VLM prompting on the full-series image (ablation without ViT4TS screening); “VLM-Short” (TAMA) refers to VLM prompting on the rolling-windows image. Only representative segments are shown for clarity.
Additional Analyses. In the Appendix, we provide further analysis of the framework, including patch‐level embedding and anomaly map visualizations, comparisons across different VLM backbones, alternative visualizations such as spectrogram baselines, and a study of how window size and patch scale affect fine‐grained detection, among other results and discussions.
Conclusion
@@
We introduce a novel two‐stage framework for unsupervised time‐series anomaly detection that reframes 1-D time series as 2-D visual inputs and leverages pretrained vision and vision–language models without any in‐domain training. In the first stage, ViT4TS applies multi‐scale cross‐patch comparisons to generate anomaly candidates in high resolution, and in the second stage, VLM4TS verifies these candidate proposals via temporal understanding with long context. Our method achieves state‐of‐the‐art performance on diverse benchmarks while using many fewer tokens.
Acknowledgments
We thank Dr. Runze Li for valuable discussions and suggestions that contributed to the methodological formulation.
Appendix A Experimental Details and Further Discussion
Benchmark Datasets
@@ Appendix A Experimental Details and Further Discussion
We use 11 datasets spanning various domains to evaluate the models’ generalizability and adaptability. The National Aeronautics and Space Administration (NASA) provided two spacecraft telemetry datasets  [^1] NASA data: https://github.com/khundman/telemanom/: Soil Moisture Active Passage (SMAP) and Mars Science Laboratory (MSL) acquired from a satellite and a rover, respectively (hundman2018lstmdt). Each numeric measurement in the target channel is accompanied by one-hot encoded information about commands sent or received by specific spacecraft modules in a given time window. The Yahoo Webscope Program provided the S5 datasets  [^2] YAHOO data: https://webscope.sandbox.yahoo.com/catalog.php?datatype=s&did=70 consisting of one set of real production traffic to Yahoo properties (A1) and three synthetic datasets (A2, A3, A4) with varying trends, noise, and pre-specified or random seasonality. The A2 and A3 datasets only contain outliers inserted at random positions, while A4 has outliers and change points. The Numenta Anomaly Benchmark (NAB) provided several datasets  [^3] NAB data: https://github.com/numenta/NAB from various domains: artificialWithAnomaly (Art), realAdExchange (AdEx), realAWSCloudwatch (AWS), realTraffic (Traffic), realTweets (Tweets). NAB is challenging due to its variety and the requirement for detectors to work across domains. Results on NAB can assess how well our approach handles completely different kinds of time series in one unified model. Similar to (wong2022aer), Table 4 summarizes the basic information of each dataset. It differentiates between real and synthetic datasets and provides the number of anomalies for each dataset. Each anomaly is classified as either point or collective, depending on the length of the anomaly. Lastly, the total number of anomalous and overall data points are provided for each dataset.
@@
Model Setup
For the primary experiment, ViT4TS uses the OpenCLIP (ilharco_gabriel_2021_5143773) implementation of CLIP with the ViT-B/16 backbone, pretrained on the LAION-400M dataset (schuhmann2021laion). For multi-scale embedding extraction, we perform average pooling over the patch grid using kernel sizes of $2\times 2$ and $3\times 3$. For the verification stage, VLM4TS employs GPT-4o (API version 2024-08-06) via OpenAI’s API.
Baseline Methods and Evaluation Procedure
@@
ARIMA (pena2013arima): autoregressive integrated moving average forecasting model; learns autocorrelations to predict future values and uses point‐wise forecast errors as anomaly scores.
@@
LSTM‐DT (hundman2018lstmdt): two‐layer LSTM for one‐step ahead prediction; uses residuals as anomaly scores.
@@
LSTM‐AE (malhotra2016lstmae): LSTM autoencoder over sliding windows; use reconstruction error as anomaly scores.
@@
VAE (park2018vae): variational autoencoder with LSTM encoder/decoder; use reconstruction likelihood and point‐wise error to construct anomaly scores.
@@
AER (wong2022aer): joint autoencoder and bidirectional LSTM predictor; optimizes reconstruction and forward/backward prediction, combining DTW‐based and point‐wise errors for anomaly scores.
@@
TadGAN (geiger2020tadgan): GAN with LSTM generator and critic; computes anomaly score via dynamic‐time‐warping, point‐wise, and area‐difference errors between real and generated series.
@@
ATrans (Anomaly Transformer) (xu2021anomaly): transformer with anomaly‐attention mechanism measuring association discrepancies among adjacent points as anomaly scores.
@@
UniTS (gao2024units): unified multi‐task transformer pretrained on diverse series; tokenizes forecasting, imputation, and anomaly tasks for cross-domain transfer. Using forecasting error as anomaly scores.
@@
TimesFM / TimesFM2 (das2024decoder): decoder‐style attention model pretrained on patched time-series corpus; using forecasting error as anomaly scores.
@@
SigLLM‐P (SigLLM‐PG on GPT, SigLLM‐PM on Mistral) (alnegheimish2024can): prompt‐based detectors that ask an LLM to mark anomalous indices in textual form.
@@
SigLLM‐D (alnegheimish2024can): detects anomalies via LLM forecasting errors—comparing next‐step predictions to actual values.
@@
TAMA (zhuang2024see): rolling‐window VLM prompts, rendering each window as an image and querying GPT-4o for anomaly detection, using confidence score as anomaly scores.
@@
All baselines except SigLLM and TAMA are evaluated using the Orion pipeline  [^4] Orion: https://github.com/sintel-dev/Orion. SigLLM is run via its repository  [^5] SIGLLM: https://github.com/sintel-dev/sigllm, while TAMA—which lacks a public implementation—is reproduced using the original prompts. TAMA is implemented on 20% of the series, due to high API costs.
Data Preprocessing, Postprocessing and Evaluation.
@@
Each raw series $x_{1:T}$ is first min–max normalized to $[0,1]$ and detrended by removing a least‐squares linear fit; for the A4 dataset, we additionally standardize before and after the known change point. After generating anomaly scores, we apply an exponentially weighted moving average (EWMA) to smooth each score (alnegheimish2022sintel; alnegheimish2022orion; hundman2018lstmdt). We then apply Gaussian quantile thresholds
$\tau_{\alpha}\;=\;\mu+k\sigma\;=\;\mu+z_{\alpha}\,\sigma,$
where $\mu$ and $\sigma$ are the mean and standard deviation and $z_{\alpha}$ is the standard normal deviate at quantile $\alpha\in\{0.10,0.01,0.001\}$, and report the unweighted contextual F1 and the maximum F1 ($\mathrm{F1\text{-}max}$) over these $z_{\alpha}$ values (hundman2018lstmdt; wong2022aer; geiger2020tadgan). Formally, let the ground-truth anomaly intervals be
$\mathbf{A}=\{\bigl(t_{s},\;t_{e}\bigr)^{i}\}_{i=1}^{m},\quad\hat{\mathbf{A}}=\{\bigl(\hat{t}_{s},\;\hat{t}_{e}\bigr)^{j}\}_{j=1}^{n}.$
Define
$\mathrm{TP}=\left|\bigl\{\,j:\exists\,i,\;\bigl(\hat{t}_{s},\hat{t}_{e}\bigr)^{j}\cap\bigl(t_{s},t_{e}\bigr)^{i}\neq\emptyset\bigl\}\right|,$
$\mathrm{FP}=\left|\bigl\{\,j:\forall\,i,\;\bigl(\hat{t}_{s},\hat{t}_{e}\bigr)^{j}\cap\bigl(t_{s},t_{e}\bigr)^{i}=\emptyset\bigl\}\right|,$
$\mathrm{FN}=\left|\bigl\{\,i:\forall\,j,\;\bigl(t_{s},t_{e}\bigr)^{i}\cap\bigl(\hat{t}_{s},\hat{t}_{e}\bigr)^{j}=\emptyset\bigl\}\right|.$
The unweighted contextual precision, recall, and F1 score are then calculated corespondingly. No interval padding is applied, ensuring that the detection performance reflects each method’s raw localization capability.
Prompt Design for VLM4TS
@@
We prompt the VLM with the full-series line plot and indicate the intervals ViT4TS marked as anomalous. Each candidate anomaly is precisely localized through the x-axis tick marks. Our goal for the VLM is to refine the output of the ViT4TS and increase the precision scores. It directs the model to discard any intervals that align with the overall trend, add missed segments exhibiting clear spikes or level shifts, and retain the exact tick-based boundaries from ViT4TS. A three-level confidence scale (1–3) supports downstream filtering. See the attached prompting box for details.
Limitation and Future Work.
While our results demonstrate VLMs’ promise for unsupervised anomaly detection, our current prompting strategy remains simple. Future work could explore advanced in-context reasoning methods, such as chain-of-thought or retrieval-augmented prompts, to further enhance temporal inference. Additionally, we have focused on univariate series; extending to multivariate data is an interesting future direction. In the following section, we give an initial exploration of a multivariate extension, showcasing how the two-stage design can be extended to incorporate multi-channel temporal information.